OpenStudy (anonymous):

How do you solve : g^2-9g-22 ?

OpenStudy (anonymous):

- Im sorry , I cant help you .

OpenStudy (anonymous):

Factor. (g - 11)( g + 2) = 0 g - 11 = 0 g = 11 g + 2 = 0 g = -2 g = 11, -2

OpenStudy (anonymous):

Thankyou mkuehm 1O

OpenStudy (amistre64):

ax^2 + bx + c = 0 a*c = m*n m+n = b (x+m) (x+n) = 0; x=-m or x=-n

OpenStudy (amistre64):

1(-22) = -22 -22 factors to; -22,1, -22+1 = -21, not it -11,2, -11+2 = - 9, thats it (x-11)(x+2) = 0 ; x = 11 or x=-2